OpenStudy (anonymous):

Write equations for the following number pattern. x y 1 7 2 49 3 343 4 2401 5 16807

Directrix (directrix):

It looks as if we are multiplying by powers of 7.

Directrix (directrix):

7 to the first power is 7. 7 to the second power is 7*7 = 49 which can be written as 7²

OpenStudy (anonymous):

yeah I know but how do you write this in a equation?

Directrix (directrix):

If x is the power, then y = 7 ^x power or is it y = x ^ 7 power? For (2,49) is 49 = 7^2 or does 49 = 2 ^ 7 ?

Directrix (directrix):

7^2 means 7 times 7. 2^7 means 2*2*2*2*2*2*2.

Directrix (directrix):

The value of y is a power of 7. Example: 2401 = 7 ^ 4 --> In the chart, this shows as (4, 2401).

Directrix (directrix):

@smithjessica27 Take a guess.

Directrix (directrix):

y = ?

Directrix (directrix):

y = 7 ^ ?

Directrix (directrix):

Y = 7^x